OpenStudy (anonymous):

if x-2 is a factor of kx^2+kx+20 then k=?

OpenStudy (anonymous):

Example 5x + 3 x + 3 / 5 is a factor. A factor sets the equation to 0. x + 3 / 5 = 0 x = -3 / 5 Substitute the value of x into the equation. 5(-3 / 5) + 3 = -3 + 3 = 0 Another example x^2 - 2x + 1 x - 1 is a factor x - 1 = 0 x = 1 Substitute the value of x into the equation. (1)^2 - 2(1) + 1 = 1 - 2 + 1 = 0

OpenStudy (akashdeepdeb):

If it is a factor then the polynomial should be = 0 Thus k2^2+k2+20 = 0 Now find the answer! :)

OpenStudy (akashdeepdeb):

I just substitued x with 2 okay?

OpenStudy (anonymous):

i did substitute x with 2 but kept getting the wrong answer. the answer is -10/3. i am getting 1.25

OpenStudy (anonymous):

Can you show your working? I got -10 / 3

OpenStudy (anonymous):

ok 1 sec

OpenStudy (akashdeepdeb):

4k + 2k = -20? Then calculate! Good job buddy! :)

OpenStudy (anonymous):

thank you so much. I kept writing the equation wrong

OpenStudy (anonymous):

k(2)^2 + k(2) + 20 = 0 4k + 2k + 20 = 0 6k + 20 = 0 6k + 20 - 20 = -20 6k = -20 6k / 6 = -20 / 6 k = -10 / 3

OpenStudy (anonymous):

i was saying k(4)+k(2) and dividing each over to get 2.5 then i was adding the leftover 1k+1k to make 2k then dividing by 2.... that's why i got 1.25. thanks!

OpenStudy (noelgreco):

Synthetic division is a quick way